# Haoran Guan

### v1

@@ 4.2 Theoretical results of SLHC2 and general settings
$\leq$ $\gamma_{s_{1}}\cdot(1+2\gamma_{m})\cdot\sqrt{1+\epsilon_{b}}\cdot\sqrt{n}\|L\|_{2}$
$\leq$ $1.1\cdot\sqrt{1+\epsilon_{b}}\cdot s_{1}{\bf u}\cdot\sqrt{n}\|L\|_{2}$
with probability at least $1-p_{f}$. (111) is proved.
Lemma 18 (Estimation of $\|E_{HQss}\|_{2}$)
For $\|E_{HQss}\|_{2}$, we can have
$\|E_{HQss}\|_{2}\leq 1.1s_{2}n^{2}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot\|L\|_{2}$ (118)
@@
Proof
For $\|E_{HQss}\|_{2}$, with Lemma 3, based on (87), we can have
$\|E_{HQss_{j}}\|_{2}\leq\gamma_{s_{2}n}\cdot\|L_{ss}\|_{2}.$ (119)
Here, $E_{HQss_{j}}$ denotes the $j$-th column of $E_{HQss}$, $j=1,2,\cdots,n$. Based on Lemma 2 and (119), we can have
$\|E_{HQss}\|_{2}$$\leq$ $\|E_{HQss}\|_{F}$ (120)
$\leq$ $\gamma_{s_{2}n}\|L_{ss}\|_{F}$
$\leq$ $1.02s_{2}n\sqrt{n}{\bf u}\|L_{ss}\|_{2}.$
Regarding $\|L_{ss}\|_{2}$, With Lemma 8, (86), (101) and (111), we can get
$\|L_{ss}\|_{2}$$\leq$ $\|\Omega_{2}\Omega_{1}L\|_{2}+\|E_{Lss}\|_{2}$ (121)
$\leq$ $\sqrt{1+\epsilon_{b}}\cdot\sqrt{n}\|L\|_{2}+1.1s_{1}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot\sqrt{n}\|L\|_{2}$
$\leq$ $1.02\sqrt{1+\epsilon_{b}}\cdot\sqrt{n}\|L\|_{2}$
with probability at least $1-p_{f}$. We put (121) into (120) and we can get
$\|E_{HQss}\|_{2}$$\leq$ $1.02s_{2}n\sqrt{n}{\bf u}\|L_{ss}\|_{2}$
$\leq$ $1.1s_{2}n^{2}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot\|L\|_{2}$
with probability at least $1-p_{f}$. (118) is proved.
Lemma 19 (Estimation of $\|G_{ss}\|_{2}$)
For $\|G_{ss}\|_{2}$, we can have
$\|G_{ss}\|_{2}\leq 1.02\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\sqrt{n}\|L\|_{2}$ (122)
@@
Proof
Regarding $\|G_{ss}\|_{2}$, when $W_{ss}$ is orthogonal, with (87), we can get
$\|G_{ss}\|_{2}$ $=$ $\|W_{ss}G_{ss}\|_{2}$ (123)
@@
$\leq$ $1.02\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\sqrt{n}\|L\|_{2}$
with probability at least $1-p_{f}$. (122) is proved.
Lemma 20 (Estimation of $\|E_{Rss}\|_{2}$)
For $\|E_{Rss}\|_{2}$, we can have
$\|E_{Rss}\|_{2}\leq 1.13n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}$ (124)
@@
Proof
To estimate $\|E_{Rss}\|_{2}$, based on Lemma 2 and (88), we can get
$\|E_{Rss}\|_{2}\leq 1.1n^{2}{\bf u}\|G_{ss}\|_{2}\|U\|_{2}.$ (125)
We put (122) into (125) and we can get
$\|E_{Rss}\|_{2}$$\leq$ $1.1n^{2}{\bf u}\|G_{ss}\|_{2}\|U\|_{2}$
$\leq$ $1.13n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}$
with probability at least $1-p_{f}$. (124) is proved.
Lemma 21 (Estimation of $\|R_{ss}^{-1}\|_{2}$)
For $\|R_{ss}^{-1}\|_{2}$, we can have
$\|R_{ss}^{-1}\|_{2}\leq\frac{1.12}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)\sigma_{min}(U)}$ (126)
@@
Proof
Regarding $\|R_{ss}^{-1}\|_{2}$, with (88), we can have
$\|R_{ss}\|_{2}$ $\leq\|G_{ss}\|_{2}\|U\|_{2}+\|E_{Rss}\|_{2}.$ (127)
We put (122) and (124) into (127) and with (102), we can have
$\|R_{ss}\|_{2}$$\leq$ $\|G_{ss}\|_{2}\|U\|_{2}+\|E_{Rss}\|_{2}$ (128)
$\leq$ $1.02\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\sqrt{n}\|L\|_{2}\|U\|_{2}$
$+$ $1.13n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}$
$\leq$ $1.04\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\sqrt{n}\|L\|_{2}\|U\|_{2}$
with probability at least $1-p_{f}$. With Lemma 4 and (88), we can have
$\sigma_{min}(R_{ss})$ $=$ $\sigma_{min}(G_{ss}U-E_{Rss})$ (129)
@@
For $\sigma_{min}(G_{ss})$, when $W_{ss}$ is orthogonal, based on Lemma 4 and (87), we can have
$\sigma_{min}(G_{ss})$ $=$ $\sigma_{min}(W_{ss}G_{ss})$ (130)
@@
For $\sigma_{min}(L_{ss})$, based on Lemma 4 and (86), we can have
$\sigma_{min}(L_{ss})\geq\sigma_{min}(\Omega_{2}\Omega_{1}L)-\|E_{Lss}\|_{2}.$ (131)
Based on (14) and (86), we can have
$\sigma_{min}(\Omega_{2}\Omega_{1}L)\geq\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)$ (132)
with probability at least $1-p_{f}$. We put (111) and (132) into (131) and we can have
$\sigma_{min}(L_{ss})\geq\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)-1.1\cdot\sqrt{1+\epsilon_{b}}\cdot s_{1}{\bf u}\cdot\sqrt{n}\|L\|_{2}$ (133)
with probability at least $1-p_{f}$. We put (133) into (130) and based on (103), we can have
$\sigma_{min}(G_{ss})$$\geq$ $\sigma_{min}(L_{ss})-\|E_{HQss}\|_{2}$ (134)
$\leq$ $\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)-1.1\cdot\sqrt{1+\epsilon_{b}}\cdot s_{1}{\bf u}\cdot\sqrt{n}\|L\|_{2}$
-- $1.1s_{2}n^{2}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot\|L\|_{2}$
@@
with probability at least $1-p_{f}$. We put (124) and (134) into (129) and based on (104), we can have
$\sigma_{min}(R_{ss})$$\geq$ $\sigma_{min}(G_{ss})\sigma_{min}(U)-\|E_{Rss}\|_{2}$ (135)
$\geq$ $0.95\cdot\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)\sigma_{min}(U)$
-- $1.13n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}$
@@
with probability at least $1-p_{f}$. With (135), we can estimate $\|R_{ss}^{-1}\|_{2}$ as
@@
with probability at least $1-p_{f}$. (126) is proved.
Lemma 22 (Estimation of $\|UR_{ss}^{-1}\|_{2}$)
For $\|UR_{ss}^{-1}\|_{2}$, we can have
$\|UR_{ss}^{-1}\|_{2}\leq\frac{1.12}{\sigma_{min}(L)}\cdot\sqrt{1-\epsilon_{s}}$ (136)
@@
Proof
To estimate $\|UR_{ss}^{-1}\|_{2}$, with (88), we can have
$\|UR_{ss}^{-1}\|_{2}\leq\|G^{-1}\|_{2}(1+\|E_{R_{ss}}\|_{2}\|R^{-1}\|_{2}).$ (137)
Regarding $\|G_{ss}^{-1}\|_{2}$, with (134), we can have
$\|G^{-1}\|_{2}$ $=$ $\frac{1}{\sigma_{min}(G_{ss})}$ (138)
$\leq$ $\frac{1.06}{\sigma_{min}(L)}\cdot\sqrt{1-\epsilon_{s}}$
with probability at least $1-p_{f}$. Therefore, we put (124), (126) and (138) into (137) and based on (104), we can have
$\|UR_{ss}^{-1}\|_{2}$$\leq$ $\|G_{ss}^{-1}\|_{2}(1+\|E_{Rss}\|_{2}\|R^{-1}\|_{2})$
$\leq$ $\frac{1.06}{\sigma_{min}(L)}\cdot\sqrt{1-\epsilon_{s}}\cdot(1+(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}\cdot\frac{1.25}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)\sigma_{min}(U)})$
$\leq$ $\frac{1.12}{\sigma_{min}(L)}\cdot\sqrt{1-\epsilon_{s}}$
with probability at least $1-p_{f}$. (136) is proved.
Lemma 23 (Estimation of $\|E_{QRss}\|_{2}$)
For $\|E_{QRss}\|_{2}$, we can have
$\|E_{QRss}\|_{2}\leq 1.15n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}\|Q_{ss}\|_{2}$ (139)
@@
Proof
To estimate $\|E_{QRss}\|_{2}$, Based on (89), with Lemma 2, we can have
$\|E_{QRss}\|_{2}$$\leq$ $1.1n^{2}{\bf u}\|Q_{ss}\|_{2}\|R_{ss}\|_{2}.$ (140)
Regarding $\|R_{ss}\|_{2}$, with (88), we can have
$\|R_{ss}\|_{2}\leq\|G_{ss}\|_{2}\|U\|_{2}+\|E_{R_{ss}}\|_{2}.$ (141)
We put (122) and (124) into (141) and based on (102), we can have
$\|R_{ss}\|_{2}$$\leq$ $\|G_{ss}\|_{2}\|U\|_{2}+\|E_{R_{ss}}\|_{2}$ (142)
$\leq$ $1.02\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\sqrt{n}\|L\|_{2}\cdot\|U\|_{2}$
$+$ $1.13n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}$
$\leq$ $1.04\sqrt{n}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}$
with probability at least $1-p_{f}$. We put (142) into (140) and we can have
$\|E_{QRss}\|_{2}$$\leq$ $1.1n^{2}{\bf u}\|Q_{ss}\|_{2}\|R_{ss}\|_{2}$
$\leq$ $1.15n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}\|Q_{ss}\|_{2}$
with probability at least $1-p_{f}$. (139) is proved.
4.4 Proof of Theorem Theorem 4.1
With the theoretical results in the previous part, we can prove Theorem 4.1 in this part. Some of the ideas are similar to that of Householder .
Proof

### v2

@@ 4.2. Theoretical results of SLHC2 and general settings
$\leq$ $1.1s_{1}{\bf u}\cdot\sqrt{s_{2}}\|\Omega_{2}\|_{2}\cdot(\sqrt{1+\epsilon_{1}}\cdot\sqrt{n}\|L\|_{2}+1.1m{\bf u}\cdot\sqrt{s_{1}n}\|\Omega_{1}\|_{2}\|L\|_{2})$
$+$ $1.1m{\bf u}\cdot\sqrt{s_{1}n}\|\Omega_{1}\|_{2}\|\Omega_{2}\|_{2}\|L\|_{2}$
$\leq$ $[1.1s_{1}{\bf u}\cdot\sqrt{1+\epsilon_{1}}\cdot\sqrt{s_{2}n}\|\Omega_{2}\|_{2}+(1+1.1s_{1}{\bf u}\cdot\sqrt{s_{2}})\cdot 1.1m{\bf u}\cdot\sqrt{s_{1}n}\|\Omega_{1}\|_{2}\|\Omega_{2}\|_{2}]\cdot\|L\|_{2}$
$\leq$ $(1.1s_{1}{\bf u}\cdot\sqrt{1+\epsilon_{1}}\cdot\sqrt{s_{2}n}\|\Omega_{2}\|_{2}+1.2m{\bf u}\cdot\sqrt{s_{1}n}\|\Omega_{1}\|_{2}\|\Omega_{2}\|_{2})\cdot\|L\|_{2}$
$=$ $t\|L\|_{2}$
with probability at least $1-p_{f}$. (4.24) is proved. ∎
Lemma 4.4 (Estimation of $\|E_{HQss}\|_{2}$).
For $\|E_{HQss}\|_{2}$, we have
(4.31) $\|E_{HQss}\|_{2}\leq 1.1s_{2}n\sqrt{n}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}$
@@
Proof.
For $\|E_{HQss}\|_{2}$, with Lemma 2.3, based on (4.3), we can have
(4.32) $\|E_{HQss_{j}}\|_{2}$$\leq$ $\gamma_{s_{2}n}\cdot\|L_{ss}\|_{2}$
$\leq$ $1.1s_{2}n{\bf u}\cdot\|L_{ss}\|_{2}.$
Here, $E_{HQss_{j}}$ denotes the $j$-th column of $E_{HQss}$, $j=1,2,\cdots,n$. Based on Lemma 2.2 and (4.32), we can have
(4.33) $\|E_{HQss}\|_{2}$$\leq$ $\|E_{HQss}\|_{F}$
$\leq$ $1.1s_{2}n{\bf u}\cdot\|L_{ss}\|_{F}$
$\leq$ $1.1s_{2}n\sqrt{n}{\bf u}\cdot\|L_{ss}\|_{2}.$
Regarding $\|L_{ss}\|_{2}$, With Lemma 2.10, (4.2), (4.11) and (4.24), we can get
(4.34) $\|L_{ss}\|_{2}$$\leq$ $\|\Omega_{2}\Omega_{1}L\|_{2}+\|E_{Lss}\|_{2}$
$\leq$ $\sqrt{1+\epsilon_{b}}\cdot\|L\|_{2}+t\|L\|_{2}$
$\leq$ $(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}$
with probability at least $1-p_{f}$. We put (4.34) into (4.33) and we can get
$\|E_{HQss}\|_{2}$$\leq$ $1.1s_{2}n\sqrt{n}{\bf u}\cdot\|L_{ss}\|_{2}$
$\leq$ $1.1s_{2}n\sqrt{n}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}$
with probability at least $1-p_{f}$. (4.31) is proved. ∎
Lemma 4.5 (Estimation of $\|G_{ss}\|_{2}$).
For $\|G_{ss}\|_{2}$, we have
(4.35) $\|G_{ss}\|_{2}\leq 1.02(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}$
@@
Proof.
Regarding $\|G_{ss}\|_{2}$, when $W_{ss}$ is orthogonal, with (4.3), we can get
(4.36) $\|G_{ss}\|_{2}$ $=$ $\|W_{ss}G_{ss}\|_{2}$
@@
$\leq$ $(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}+1.1s_{2}n\sqrt{n}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}$
$\leq$ $(1+1.1s_{2}n\sqrt{n}{\bf u})\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}$
$\leq$ $1.02\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}$
with probability at least $1-p_{f}$. (4.35) is proved. ∎
Lemma 4.6 (Estimation of $\|E_{Rss}\|_{2}$).
For $\|E_{Rss}\|_{2}$, we have
(4.37) $\|E_{Rss}\|_{2}\leq 1.13n^{2}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}\|U\|_{2}$
@@
Proof.
To estimate $\|E_{Rss}\|_{2}$, based on Lemma 2.2 and (4.4), we can get
(4.38) $\|E_{Rss}\|_{2}\leq 1.1n^{2}{\bf u}\cdot\|G_{ss}\|_{2}\|U\|_{2}.$
We put (4.35) into (4.38) and we can get
$\|E_{Rss}\|_{2}$$\leq$ $1.1n^{2}{\bf u}\cdot\|G_{ss}\|_{2}\|U\|_{2}$
$\leq$ $1.13n^{2}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}\|U\|_{2}$
with probability at least $1-p_{f}$. (4.37) is proved. ∎
Lemma 4.7 (Estimation of $\|R_{ss}^{-1}\|_{2}$).
For $\|R_{ss}^{-1}\|_{2}$, we have
(4.39) $\|R_{ss}^{-1}\|_{2}\leq\frac{1.12}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)\sigma_{min}(U)}$
@@
Proof.
Regarding $\|R_{ss}^{-1}\|_{2}$, with (4.4), we can have
(4.40) $\|R_{ss}\|_{2}$ $\leq\|G_{ss}\|_{2}\|U\|_{2}+\|E_{Rss}\|_{2}.$
We put (4.35) and (4.37) into (4.40) and with (4.12), we can have
(4.41) $\|R_{ss}\|_{2}$$\leq$ $\|G_{ss}\|_{2}\|U\|_{2}+\|E_{Rss}\|_{2}$
$\leq$ $1.02\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}\|U\|_{2}$
$+$ $1.13n^{2}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}\|U\|_{2}$
with probability at least $1-p_{f}$. With Lemma 2.5 and (4.4), we can have
(4.42) $\sigma_{min}(R_{ss})$ $=$ $\sigma_{min}(G_{ss}U-E_{Rss})$
@@
For $\sigma_{min}(G_{ss})$, when $W_{ss}$ is orthogonal, based on Lemma 2.5 and (4.3), we can have
(4.43) $\sigma_{min}(G_{ss})$ $=$ $\sigma_{min}(W_{ss}G_{ss})$
@@
For $\sigma_{min}(L_{ss})$, based on Lemma 2.5 and (4.2), we can have
(4.44) $\sigma_{min}(L_{ss})\geq\sigma_{min}(\Omega_{2}\Omega_{1}L)-\|E_{Lss}\|_{2}.$
Based on (2.14) and (4.2), we can have
(4.45) $\sigma_{min}(\Omega_{2}\Omega_{1}L)\geq\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)$
with probability at least $1-p_{f}$. We put (4.24) and (4.45) into (4.44) and we can have
(4.46) $\sigma_{min}(L_{ss})\geq\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)-t\|L\|_{2}$
with probability at least $1-p_{f}$. We put (4.46) into (4.43) and based on (4.13), we can have
(4.47) $\sigma_{min}(G_{ss})$$\geq$ $\sigma_{min}(L_{ss})-\|E_{HQss}\|_{2}$
$\geq$ $\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)-t\|L\|_{2}-1.1s_{2}n\sqrt{n}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}$
$\geq$ $\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)-[(1+1.1s_{2}n\sqrt{n}{\bf u})\cdot t+1.1s_{2}n\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}]\cdot\|L\|_{2}$
$\geq$ $\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)-(1.02t+1.1s_{2}n\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}})\cdot\|L\|_{2}$
@@
with probability at least $1-p_{f}$. We put (4.37) and (4.47) into (4.42) and based on (4.14), we can have
(4.48) $\sigma_{min}(R_{ss})$$\geq$ $\sigma_{min}(G_{ss})\sigma_{min}(U)-\|E_{Rss}\|_{2}$
$\geq$ $0.95\cdot\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)\sigma_{min}(U)-1.13n^{2}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}\|U\|_{2}$
@@
with probability at least $1-p_{f}$. With (4.48), we can estimate $\|R_{ss}^{-1}\|_{2}$ as
@@
with probability at least $1-p_{f}$. (4.39) is proved. ∎
Lemma 4.8 (Estimation of $\|UR_{ss}^{-1}\|_{2}$).
For $\|UR_{ss}^{-1}\|_{2}$, we have
(4.49) $\|UR_{ss}^{-1}\|_{2}\leq\frac{1.12}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)}$
@@
Proof.
To estimate $\|UR_{ss}^{-1}\|_{2}$, with (4.4), we can have
(4.50) $\|UR_{ss}^{-1}\|_{2}\leq\|G_{ss}^{-1}\|_{2}(1+\|E_{R_{ss}}\|_{2}\|R^{-1}\|_{2}).$
Regarding $\|G_{ss}^{-1}\|_{2}$, with (4.47), we can have
(4.51) $\|G_{ss}^{-1}\|_{2}$ $=$ $\frac{1}{\sigma_{min}(G_{ss})}$
$\leq$ $\frac{1.06}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)}$
with probability at least $1-p_{f}$. Therefore, we put (4.37), (4.39) and (4.51) into (4.50) and based on (4.14), we can have
$\|UR_{ss}^{-1}\|_{2}$$\leq$ $\|G_{ss}^{-1}\|_{2}(1+\|E_{Rss}\|_{2}\|R_{ss}^{-1}\|_{2})$
$\leq$ $\frac{1.06}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)}\cdot[1+1.13n^{2}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}\|U\|_{2}\cdot\frac{1.12}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)\sigma_{min}(U)}]$
$\leq$ $\frac{1.12}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)}$
with probability at least $1-p_{f}$. (4.49) is proved. ∎
Lemma 4.9 (Estimation of $\|E_{QRss}\|_{2}$).
For $\|E_{QRss}\|_{2}$, we have
(4.52) $\|E_{QRss}\|_{2}\leq 1.15n^{2}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}\|U\|_{2}\|Q_{ss}\|_{2}$
@@
Proof.
To estimate $\|E_{QRss}\|_{2}$, Based on (4.5), with Lemma 2.2, we can have
(4.53) $\|E_{QRss}\|_{2}$$\leq$ $1.1n^{2}{\bf u}\cdot\|Q_{ss}\|_{2}\|R_{ss}\|_{2}.$
Regarding $\|R_{ss}\|_{2}$, with (4.4), we can have
(4.54) $\|R_{ss}\|_{2}\leq\|G_{ss}\|_{2}\|U\|_{2}+\|E_{R_{ss}}\|_{2}.$
We put (4.35) and (4.37) into (4.54) and based on (4.12), we can have
(4.55) $\|R_{ss}\|_{2}$$\leq$ $\|G_{ss}\|_{2}\|U\|_{2}+\|E_{R_{ss}}\|_{2}$
$\leq$ $1.02\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}\|U\|_{2}$
$+$ $1.13n^{2}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}\|U\|_{2}$
$\leq$ $1.04\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}\|U\|_{2}$
with probability at least $1-p_{f}$. We put (4.55) into (4.53) and we can have
$\|E_{QRss}\|_{2}$$\leq$ $1.1n^{2}{\bf u}\cdot\|Q_{ss}\|_{2}\|R_{ss}\|_{2}$
$\leq$ $1.15n^{2}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}\|U\|_{2}\|Q_{ss}\|_{2}$
with probability at least $1-p_{f}$. (4.52) is proved. ∎
4.4. Proof of Theorem Theorem 4.1
With the theoretical results in the previous part, we can prove Theorem 4.1 in this part. Some of the ideas are similar to those of [2].
Proof.